What annual payment would discharge a debt of rs.1025 due in 2 years at the rate of 5% compound interest?

Grade:8

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
We know in compound interest formula for present worth of Rs. x due in T years is:
PW=x/(1+ R/100)^T
So here, present worth od x due 1 year+ present worth of x due 2 years=1025
putting formula we get,
x/(1+ 5/100)^1 + x/(1+ 5/100)^2 =1025
solving, we get
x= 205*441 / 164
=>x=Rs. 551.25
So, the annual payment required is Rs. 551.25 . [Ans]
Thanks,
Soumya Ranjan Mohanty.